### v1

# Prediction of Protein Three-dimensional Structures via a Hardware-Executable Quantum Computing Framework

Abstract
Accurate prediction of protein active site structures remains a central challenge in structural biology, particularly for short and flexible peptide fragments where conventional methods often fail. Here, we present a quantum computing framework specifically developed for utility-level quantum processors to address this problem. Starting from an amino acid sequence, we formulate the structure prediction task as a ground-state energy minimization problem using the Variational Quantum Eigensolver (VQE). Amino acid connectivity is encoded on a tetrahedral lattice model, and structural constraints—including steric, geometric, and chirality terms—are mapped into a problem-specific Hamiltonian expressed as sparse Pauli operators. The optimization is executed via a two-stage architecture separating energy estimation and measurement decoding, allowing noise mitigation under realistic quantum device conditions. We evaluate the framework on 23 randomly selected real protein fragments from the PDBbind dataset, as well as 7 real fragments from proteins with therapeutic potential, and run the experiments on the IBM–Cleveland Clinic quantum processor. Structural predictions are benchmarked against AlphaFold3 (AF3) using identical postprocessing and docking procedures. Our quantum method outperformed AF3 in both RMSD (Root-Mean-Square Deviation) and docking efficacy. This work demonstrates, for the first time, a complete end-to-end pipeline for biologically relevant structure prediction on real quantum hardware, highlighting its engineering feasibility and practical advantage over existing classical and deep learning approaches.
Introduction
Predicting the local three-dimensional structures of proteins, particularly at ligand-binding sites, is a central problem in structural biology with broad implications for molecular function annotation, drug design, and protein engineering[1]. The conformational landscape of these regions is often characterized by high flexibility and functional specificity, making accurately model structures both scientifically critical and technically challenging[2]. Traditional physics-based methods, such as molecular dynamics (MD), can in principle sample native-like structures but suffer from prohibitive computational costs and scalability issues[3, 4]. On the other hand, recent breakthroughs in deep learning, such as AlphaFold2 and AlphaFold3, have demonstrated remarkable success in full-length protein structure prediction[5, 6]. However, their performance substantially deteriorates when applied to short peptide fragments or highly flexible local domains, where limited sequence context and sparse evolutionary information lead to structural ambiguities and reduced accuracy [7]. In particular, short protein fragments involved in ligand binding often contain fewer than 20 amino acids and exhibit complex energetic constraints that are difficult to capture with purely data-driven models. Most of the results predicted by deep learning methods are excellent; however, in some cases, issues such as overfitting or incorrect predictions may occur. Deep learning approaches still have certain blind spots, particularly in the prediction of short fragments [8]. Although several methods (e.g. modeling the problem as a binary optimization task for optimization, employing quantum annealing algorithms, or using the QAOA (Quantum Approximate Optimization Algorithm) method) have previously been proposed for protein structure prediction using quantum computing, their validation in real-world scenarios remains highly limited. Some approaches are only capable of generating predictions, but the performance of these predictions often falls short of being practically applicable to real problems (e.g. the docking application discussed in this paper) [9]. Therefore, there is an increasing need for protein prediction frameworks that operate under limited input information capacity (e.g., short protein fragments), make efficient use of quantum resources, and deliver high-precision results to meet the requirements of tasks such as docking. Quantum computing offers a new paradigm for addressing classically intractable optimization problems by exploiting quantum parallelism, entanglement, and interference. While current quantum hardware remains limited by noise, decoherence, and qubit count—commonly referred to as utility-level processors—it has reached a level of maturity that enables real-world applications in small-scale, high-precision tasks. Among hybrid quantum-classical algorithms, the Variational Quantum Eigensolver (VQE) is particularly well-suited to molecular energy landscape optimization [10]. By encoding molecular constraints into parameterized quantum circuits, VQE enables the estimation of ground-state energies for quantum Hamiltonians constructed from physical models.
In this work, we present an end-to-end quantum computing framework for predicting the local structure of protein binding site fragments using utility-level quantum hardware. Our approach begins with an amino acid sequence and maps its spatial configuration to a tetrahedral lattice, where each residue is represented as a node with four directional connectivity options. These configurations are encoded as quantum measurement states corresponding to discrete Hamiltonians that incorporate steric constraints, chirality terms, geometric restrictions, and amino acid interactions. To mitigate the effects of hardware noise, we design a two-stage execution architecture that separates energy estimation from structural decoding. This strategy enhances execution stability while maintaining fidelity in structural recovery. We validate our framework using 30 short peptide sequences extracted from the PDBbind database [11, 12, 13]. The sequences are processed through our quantum pipeline and executed on the IBM-Cleveland Clinic 127-qubit superconducting quantum processor. For benchmarking, we compare the resulting structures against those predicted by AlphaFold3, followed by uniform postprocessing and ligand docking using AutoDock[14, 15, 16, 17, 6]. Our quantum method outperforms AlphaFold3 in both docking success rate and RMSD (Root-Mean-Square Deviation), demonstrating the feasibility and effectiveness of quantum structure prediction on real quantum hardware. This study marks one of the first demonstrations of quantum computing applied to biologically relevant structure prediction on real devices. It establishes a blueprint for leveraging utility-level quantum processors in solving domain-specific, engineering-scale problems that remain out of reach for both classical simulations and data-driven models.
Results
A Quantum Engineering Framework for Structure Prediction
To enable the prediction of protein binding site fragment structures on real quantum hardware, we developed a fully deployable, end-to-end quantum computing framework tailored for utility-level quantum processors (Figure 1). This framework takes amino acid sequences as input and outputs three-dimensional conformations that satisfy physical and geometric constraints. The system is designed with engineering completeness, modular extensibility, and hardware compatibility in mind, supporting downstream structural and functional evaluations.
The framework consists of the following core stages: (1) Sequence-based quantum modeling, where the input amino acid sequence is used to construct a quantum formulation of the structural problem, incorporating Miyazawa–Jernigan interactions and additional geometric constraints, which are encoded as Pauli operators. (2) Parameterized circuit construction, where the number of required qubits is determined and a parameterized quantum circuit is built using a selected ansatz, with random initialization of parameters and Hamiltonian association. (3) Hybrid variational optimization, in which the Qiskit Runtime middleware orchestrates VQE execution across quantum and classical resources, iteratively updating circuit parameters to minimize system energy on real quantum hardware. (4) Measurement and structural decoding: A new fixed-parameter quantum circuit is constructed by combining the optimized parameters with the same ansatz, followed by a second-stage compilation to produce a hardware-executable measurement circuit. The circuit is then executed with a fixed number of shots as specified by the user-defined precision. The resulting bitstrings are statistically processed, and the most probable output is reverse-mapped into a spatial backbone vector. And (5) Postprocessing for downstream usability, including atom completion and charge neutralization to generate biologically usable protein structures for tasks such as docking or simulation. As illustrated in Figure 1, the framework takes an amino acid sequence represented as a string as input. This sequence is first modeled using the Miyazawa–Jernigan interaction [18] along with additional structural constraints. The modeled problem is then translated into the quantum domain in the form of Pauli operators. These Pauli operators are used to determine the required number of qubits and to construct the Hamiltonian of the system. Based on the Pauli formulation, the algorithm automatically constructs a parameterized quantum circuit using a selected ansatz and initializes the variational parameters randomly. The execution is managed through a Qiskit Runtime session [19], which handles job orchestration. The parameterized circuit and Hamiltonian are uploaded to the quantum processor, where a second-stage compilation is performed. A cross-platform hybrid Variational Quantum Eigensolver (VQE) is then executed, combining quantum sampling with a classical optimizer running locally. The optimization loop in the VQE continues iterating until the minimum energy is found or a predefined termination condition is satisfied. At convergence, the optimal set(s) of variational parameters are bound to the predefined ansatz, and the circuit is recompiled. The finalized quantum circuit is then executed multiple times according to a user-defined precision to perform quantum state measurements. Upon completion of the measurement task, results are retrieved. The measured outcomes are statistically processed on the classical backend, and the highest-probability result—represented as a bitstring (binary string)—is selected. This bitstring is then reverse-mapped using the same encoding rules defined during problem modeling, reconstructing the predicted protein backbone structure. A final postprocessing stage is performed locally, including structural refinement steps such as atom completion and charge neutralization, in preparation for downstream tasks such as molecular docking.
Each residue in the input sequence is mapped to a node on a tetrahedral grid, where four directional edges represent possible backbone extensions. The spatial connectivity between adjacent residues is encoded into segments of a quantum bitstring, which defines the geometric path of the chain. To model physical validity, we construct a sparse Pauli Hamiltonian incorporating chirality constraints, steric repulsion, geometric feasibility, and amino acid interaction energies. This Hamiltonian is optimized using VQE to obtain a quantum state corresponding to the lowest-energy configuration. Given the constraints of current utility-level hardware—including short coherence times, measurement noise, and gate infidelity—we introduce a two-phase execution strategy. In the first phase, the parameterized quantum circuit is used for energy estimation and variational optimization. In the second phase, the optimized parameters from the first phase are fixed and compiled into a stable circuit composed entirely of native quantum gates supported by the processor. This fixed circuit is then measured repeatedly to obtain bitstrings at a user-specified precision. The decoupling of optimization and measurement enhances robustness to noise, increases execution stability, and ensures the reproducibility of structural predictions. The final output of the framework is a set of spatially consistent three-dimensional structures that can be evaluated through structural analysis and ligand docking.
Figure 1: End-to-End Hardware-Executable Quantum Workflow for Protein Structure Prediction. The framework takes an amino acid sequence as input, encodes structural constraints into Pauli operators, and constructs a parameterized quantum circuit using a selected ansatz. A hybrid VQE is executed on IBM Quantum hardware via Qiskit Runtime to optimize circuit parameters. The lowest-energy binary outcome is reverse-mapped to a predicted backbone structure, followed by postprocessing for downstream applications. The framework consists of five submodules: (1) Problem Modeling Module: Models the input sequence as Pauli operators and constructs the corresponding Hamiltonian. (2) Quantum Circuit Construction Module: Selects the circuit architecture, builds a parameterized quantum circuit, and associates it with the Hamiltonian. (3) Hybrid Variational Optimization Module: Optimizes the parameters of the quantum circuit using a quantum-classical hybrid algorithm (VQE) to search for low-energy states. (4) Measurement Module: Reconstructs a fixed-parameter quantum circuit using the optimized parameters, performs measurements, and decodes the results. (5) Postprocessing Module: Processes the predicted outputs to generate complete protein structures suitable for downstream biological tasks such as docking or simulation.
Execution on Utility-Level Quantum Hardware
To validate the executability and engineering robustness of the proposed framework on real quantum hardware, we deployed the full structure prediction pipeline on the IBM-Cleveland Clinic’s 127-qubit superconducting quantum processor. All tasks were executed directly on physical hardware rather than in simulation environments. Depending on the fragment length and the complexity of encoded amino acid interactions, the number of active qubits required for each task ranged from several dozen to over a hundred. We used a tetrahedral lattice model to discretize the amino acid sequence and constructed a sparse Pauli Hamiltonian incorporating steric repulsion, chirality constraints, geometric feasibility, and residue-level interaction energies. During the variational optimization phase, we employed a parameterized quantum circuit composed of native single-qubit rotation gates (e.g., $\text{Ry}(\theta)$) and controlled-NOT (CX) gates. These circuits were compiled and topology-mapped using Qiskit’s transpiler to ensure compatibility with the IBM-Cleveland Clinic quantum processor. To improve tolerance to hardware noise and maintain structural prediction stability, we adopted a two-stage execution strategy, architecturally separated into a variational optimization module and a dedicated measurement module. In the first stage, the Variational Quantum Eigensolver (VQE) was used to minimize the expectation value of the constructed Hamiltonian. Each energy evaluation was performed using 2,000 shots for quantum energy estimation, combined with at least 200 iterations of the COBYLA (Constrained Optimization BY Linear Approximations) classical optimizer. In the second stage, the optimized parameter set from the first phase was fixed and used to reconstruct a non-parametric circuit, which was then recompiled into a stable quantum circuit composed entirely of native gates. This fixed circuit was executed 20,000 times to obtain measurement bitstrings representing structural states. This two-phase process ensured reproducibility and robustness under noise-constrained physical environments. For testing, we selected 23 representative peptide fragments from the PDBbind database [11, 12, 13]. Each fragment originated from a distinct protein-ligand complex with experimentally determined receptor structures. These fragments were chosen to be structurally independent and diverse in both residue composition and sequence ordering, enabling us to assess the generalization capability of our framework across varying input conditions. Table 1 summarizes the selected samples’ PDB IDs, residue sequences, sequence lengths, and execution parameters including number of qubits, circuit depth, execution time, and the system’s minimum, maximum, and energy range during optimization. Due to the high computational cost of quantum hardware, we adopted a strategy of low-iteration VQE optimization coupled with high-precision measurements. If resource constraints were relaxed, it would be feasible to achieve better convergence paths and structure prediction accuracy by increasing the number of optimization steps and adopting gradient-based optimizers. The reason why the execution times shown in the table do not increase proportionally with the length of the protein fragments is mainly due to the current task scheduling mechanism on quantum computers. Since the scheduling strategy is not yet fully intelligent, some tasks may be preempted by higher-priority jobs—such as those submitted by accounts with higher access levels or with shorter estimated runtimes—which results in additional session delays and idle time.
The 23 test cases ranged from short peptides with 5 amino acids to medium-length chains of 10–14 residues, ensuring the framework’s compatibility across varying sequence lengths. Several clear trends were observed from the collected execution data: first, the number of qubits increased significantly with sequence length, as more ancillary qubits are required to represent growing residue-residue interactions and constraints. Second, the system energy exhibited a stepwise increase as sequence length increased, with sharp jumps observed at certain threshold lengths. Using the system’s energy range as an indicator, we found that increasing the sequence length from 6 to 7 residues led to a jump in energy range from approximately 260 a.u. (atomic units, Hartree energy) to over 600 a.u., an increase of about 131%. From 7 to 9 residues, energy range increased from around 610 a.u. to approximately 1,350 a.u. (+121%), and further to over 2,100 a.u. at 10 residues (+56%). Between 10 and 11 residues, energy range continued to rise to 2,500–2,700 a.u. (+20–30%), then reached 4,600–5,300 a.u. at 13 residues (+85%), and finally surpassed 9,000 a.u. at 14 residues (+72% over 13 residues and +2,700% over the initial 6-residue case). Both the minimum and maximum energy values also expanded rapidly with sequence length. For example, minimum energy increased from approximately 10 a.u. (5 residues) to over 23,000 a.u. (14 residues), while maximum energy increased from  15 a.u. to over 32,000 a.u.—representing increases of approximately 230,000% and 210,000%, respectively. These nonlinear surges demonstrate the exponential growth of the quantum energy landscape as conformational dimensionality increases. This nonlinear energy amplification reveals two key insights. First, the quantum system is capable of encoding subtle structural variations with high precision: even the addition or removal of a single amino acid causes energy changes distinguishable by the quantum model. Second, the energy of the quantum system, as encoded under our Hamiltonian mapping, is highly correlated with the underlying conformational energy of the protein fragment. Thus, the lowest-energy quantum state produced by VQE directly corresponds to the predicted structure, validating the soundness of our prediction strategy. It is precisely due to this energy amplification behavior that we modularized the measurement stage, allowing it to operate independently with non-parametric circuits for high-precision bitstring sampling. This design preserves prediction fidelity while maintaining engineering tractability under realistic hardware constraints.
Table 1: Summary of quantum structure prediction tasks on IBM-Cleveland Clinic. Each row corresponds to a predicted protein fragment with its amino acid sequence, number of qubits, energy profile, execution time, and circuit depth.
PDB ID Sequence Len Residues Qubits Min E Max E Range Time (s) Depth
4aoi VVLPYMKHGDLRNF 14 1155–1168 102 23245.37 32378.95 9133.58 13328.65 413
4cig VRDQAEHLKTAVQM 14 165–178 102 21375.59 29846.54 8470.94 17293.54 413
4tmk IEGLEGAGKTTARN 14 8–21 102 22590.21 29135.42 6545.21 199292.66 413
3d7z YLVTHLMGADLNNI 14 103–116 102 22979.86 29707.30 6727.43 156289.48 413
4jpx DYLEAYGKGGVKA 13 154–166 92 16962.09 22231.95 5269.86 90422.62 373
1yc4 ELISNSSDALDKI 13 47–59 92 16129.38 20745.81 4616.42 15777.29 373
5cqu RKLGRGKYSEVFE 13 43–55 92 17865.39 22801.51 4936.12 7620.94 373
2qbs HCSAGIGRSGT 11 214–224 72 6691.57 9356.87 2665.30 13899.11 293
4f5y GLAWSYYIGYL 11 158–168 72 6408.50 8858.60 2450.10 6157.46 293
5cxa FDGKGGILAHA 11 174–184 72 6946.43 9298.82 2352.40 5638.71 293
4mc1 LLDTGADDTV 10 23–32 63 4092.24 6199.23 2106.99 5609.02 257
3d83 YLVTHLMGAD 10 103–112 63 4235.34 6119.16 1883.82 19833.57 257
3vf7 LLDTGADDTV 10 23–32 63 3975.02 6162.42 2187.40 5348.25 257
4y79 DACQGDSGG 9 189–197 54 1549.16 2874.21 1325.05 207445.70 221
1e2l AQITMGMPY 9 124–132 54 1509.66 2837.82 1328.15 12951.69 221
1gx8 SAPLRVYVE 9 36–44 54 1626.02 3053.53 1427.51 14080.77 221
1m7y TAGATSANE 9 117–125 54 1420.38 2714.98 1294.60 12918.04 221
3dx3 HNDPGWI 7 90–96 38 339.99 962.62 622.63 4661.24 157
5c28 CDLCSVT 7 663–669 38 386.81 792.78 405.97 114029.96 157
3ibi IQFHFH 6 91–96 23 120.66 455.42 334.76 4486.62 97
2v25 ATFTIT 6 81–86 23 100.42 340.83 240.42 22356.46 97
3ckz VKDRS 5 149–153 12 10.43 14.65 4.22 5763.36 53
3eax RYRDV 5 45–49 12 10.36 16.02 5.66 4028.72 53
Structural Evaluation and Docking Consistency
To comprehensively evaluate the practical performance of our framework in protein structure prediction tasks, we conducted a systematic analysis of the predicted structures from two perspectives: geometric accuracy and functional consistency. Geometric accuracy was assessed using the RMSD (Root-Mean-Square Deviation), which measures the structural proximity between the quantum-predicted and experimentally determined conformations. Functional consistency was evaluated based on AutoDock scoring functions, comparing the binding affinity of docking simulations using different predicted structures.
Our RMSD-based geometric comparison across 23 protein binding site fragments indicates that the quantum method achieved lower RMSD values in 18 samples, outperforming AlphaFold3 (AF3), which had better results in the remaining 5 cases (Figure S1). The average RMSD for quantum predictions was 3.33 Å, with a median of 3.53 Å, while AF3 predictions yielded an average RMSD of 3.87 Å and a median of 3.92 Å. Notably, the quantum method exhibited stronger structural fidelity in shorter sequences (5-9 residues) and in regions with higher structural flexibility. The predicted structures were used as ligands in docking experiments with a standardized receptor structure, and binding affinity were recorded over multiple runs. Results demonstrate that quantum-predicted structures achieved better binding affinities in 21 out of 23 cases (Figure S2). The average binding affinity for the quantum method was -4.38 kcal/mol, while the AF3-predicted structures averaged -4.00 kcal/mol. This suggests that quantum predictions not only exhibit superior geometric precision, but also more accurately capture the energetic characteristics of real biomolecular interactions. Furthermore, a cross-analysis of structural and functional metrics reveals that in most cases, quantum-predicted structures simultaneously exhibit lower RMSD and more favorable binding affinities, indicating a strong correlation between geometric and energetic consistency. This pattern suggests that, within the current problem scale, quantum predictions effectively capture both low-energy conformations and biologically relevant binding configurations. In summary, the quantum computing framework presented in this study delivers structural outputs on real quantum hardware that outperform state-of-the-art deep learning methods in both geometric and functional aspects. Particularly in low-information or small-sample prediction scenarios, where traditional models tend to degrade, our framework demonstrates significant advantages, highlighting its potential utility and scalability in structural biology and bioengineering applications.
Case Studies in Real Applications
[…]
Among the seven cases described above, we selected two representative examples: 6mu3 and 2xxx to visualize their quantum-system energy landscapes and to illustrate, within the practical engineering framework, the corresponding resource usage and native-gate quantum circuits. The two examples were chosen for the following reasons. First, they lie at opposite ends of the sequence-length spectrum—one peptide contains five residues and the other twelve—so their qubit requirements, circuit depths, and total energy scales differ markedly, providing a clear basis for analysing how system characteristics change with problem size. Second, the patterns of inter-residue interactions in the two sequences are entirely different, yielding substantially different Hamiltonians; this contrast highlights the versatility and general applicability of our framework. In the case of 6mu3, shown in Figure 2A-D, we selected the five lowest-energy conformations for visualization. The sequence numbers here represent the ranking starting from the lowest energy. For horizontal comparison, we conducted docking tests using the conformations corresponding to a few other low-energy states, excluding the ground-state energy conformation, with the same receptor molecule in the dataset. During testing, we performed 20 docking tests for each predicted structure, using a different random seed for each test. Each test retained 9 result sets, yielding a total of 180 possible docking outcomes. The final docking score was calculated by averaging three metrics—affinity, RLB (RMSD lower bound), and RUB (RMSD upper bound)—across all 180 results, forming a three-dimensional scoring system. In this case, the energy differences for conformations ranked 2–4 were within the range of ±0.05, and their structures were identical after reverse mapping. These energy differences were caused by perturbations in the quantum system. However, the error magnitude was within a percentage level, so it did not affect solving practical problems. The affinity of these four conformations with the same receptor molecule was greater than 0, indicating that they could not dock with the receptor molecule. The conformation corresponding to the lowest energy point successfully docked, validating the rationality of the method. In the case of 2xxx, shown in Figure 2E-G, we selected six conformations corresponding to the lowest system energies. Owing to the increased number of amino acids in the second instance, the energy span of the quantum system increased significantly, which resulted in these six conformations being distinct from one another. We applied the same testing method used in the example shown in Figure 2A-D for horizontal comparison on this set of results. In the tests, a clear positive correlation was observed between the quantum system energy of the conformations and their docking affinity. Specifically, as the quantum system energy of the conformations decreased, the resulting conformations became increasingly suitable for docking and closer to the actual docking structure.
Figure 2: Details of two example cases in real-world application scenarios. (A–D) Case1, 6mu3 (YAGYS): (A) The quantum circuit composed of native gates of the quantum processor, generated after compiling the modeled problem for Case 1. (B) The distribution plot of the measured quantum states, where the tallest bar represents the most frequently observed outcome. This highest-probability result corresponds to the final expected measurement outcome. (C) The actual qubits occupied by Case 1 on the quantum processor. (D) VQE energy trace with five lowest-energy conformers (insets: docking affinity, RLB, RUB). (E–G)Case2, 2xxx (GAVEDGATMTFF): (E) The quantum circuit composed of native gates of the quantum processor, generated after compiling the modeled problem for Case 2. (F) The actual qubits occupied by Case 2 on the quantum processor. (G) VQE energy trace with six lowest-energy conformers (insets: docking metrics).
Compared to AlphaFold3 predictions, the quantum method’s predicted structures achieved approximately 20% reductions in the upper and lower bounds of RMSD in different test cases during docking tests, indicating higher consistency and accuracy in structural prediction (Figure 3H-I). In terms of affinity score, the quantum method achieved an average improvement of about 10% over AlphaFold3 (Figure 3G). Additionally, the quantum method exhibited a tighter distribution of affinity scores, with significantly fewer outliers, further validating its reliability in optimizing docking performance. For protein structure prediction, besides evaluating predictive accuracy in docking tasks through binding affinity scores and RMSD bounds of docking poses, another important metric is the RMSD itself. The RMSD (root-mean-square deviation) metric is used to quantify structural differences between the predicted and experimental structures. In our experiments, we compared the RMSD of structures predicted by the VQE method and those predicted by AF3 against the corresponding experimental structures in the dataset. As shown in Figure 3F, the VQE method outperformed AF3 in six out of seven cases. For 6mu3, due to its short chain length and the limited degrees of freedom of the modeling approach, it lagged behind AF3 by only a percentage-level margin. Although most experimental results did not significantly surpass AF3 predictions due to limitations in modeling precision, they nevertheless demonstrate that quantum methods possess certain advantages.
Here we select a representative experimental example to further illustrate the results. As shown in Figure 3B-E, we chose a fragment from the binding pocket region of protein 3b26 for demonstration. In our experiments, this fragment consists of 11 amino acids—one of the longer test cases—which facilitates clear visualization. In the visualized outcome, the structure predicted by the quantum method clearly and accurately captures the overall fold and key structural features of the selected segment, showing a high degree of similarity to the experimentally determined structure in the dataset. In contrast, the AF3 prediction shown in Figure E deviates substantially from the experimental result: AF3 models this region as an $\alpha$-helical secondary structure, which differs markedly from the true conformation and loses essential structural characteristics. In our experiments, many AF3 predictions tended to default to helical structures for unknown regions. We hypothesize that this bias arises because helices have high recognizability and pronounced patterns, which create an information “pit” in pattern-recognition models; under limited system information capacity, the deep learning model gravitates toward predictions with the most distinctive features, thus sacrificing accuracy. In contrast, quantum methods build from first principles and do not fall into pattern-recognition traps. Even when system information capacity is severely constrained, they can still produce results with reliable fidelity.
Figure 3: Structural testing and visual representation (A) The docking workflow. The predicted structures from both methods undergo hydrogen addition and charge neutralization, followed by docking tests with the same receptor molecule from the dataset. To ensure fairness in evaluation, both prediction results are processed using an identical postprocessing pipeline. (B) Native structure of 3b26. (C) Docking pocket view. (D) Quantum overlay. The quantum method essentially captures the overall trajectory of the regional fragment and the backbone amino acid coordinates, showing high consistency with the reference (scanned) structure. (E) AF3 overlay. The prediction from AF3 appears to have fallen into a locally optimal state, exhibiting a tendency to converge toward structurally prominent features when the available information capacity (i.e., amino acid chain length) is limited. (F) RMSD comparison. Comparison of the structural differences in RMSD between the predictions of the two methods and the reference structures from the dataset. A smaller RMSD value indicates that the predicted structure is closer to the actual structure, reflecting higher prediction accuracy. (G) Affinity comparison. (H) RMSD upper bound comparison. (I) RMSD lower bound comparison.
Scalability
[…]
Discussion
This study provides the validation of the feasibility and potential of quantum computing, particularly the Variational Quantum Eigensolver (VQE), in predicting protein structures. By mapping the protein conformation problem into the Hamiltonian optimization problem of a quantum system, quantum computing demonstrates unique advantages in handling molecular energy domains. Through parameterized quantum circuits, quantum systems efficiently generate energy values and iteratively refine them using classical optimizers, rapidly converging to the lowest-energy conformation [33]. In comparison to existing AI methods such as deep learning, quantum methods offer higher theoretical reliability. While AI-based approaches, such as AlphaFold3 [6], achieve exceptional efficiency in predicting new sequences after model training, they are inherently probabilistic models reliant on large-scale data and suffer from limited interpretability [34]. These characteristics may result in limitations when addressing local structures or specific cases. In docking tasks involving small-scale fragments similar to those we used, the focus is on interface pocket surfaces formed by short peptide chains that require highly accurate structural predictions. For deep learning models like AlphaFold3 [6], although these models are built with neural networks containing millions of parameters, the information capacity for predicting such structures is extremely limited. When predicting these structures, the models are prone to falling into ”information traps,” such as generating structures that appear complete and meet basic amino acid distance characteristics but fail to satisfy amino acid residue interactions or chirality constraints. As a result, these structures do not meet the requirements of docking tasks. In contrast, quantum methods model protein conformations from the foundation of physical principles, circumventing the probabilistic constraints of AI methods and providing theoretically rigorous predictions. Additionally, quantum methods avoid the high computational costs associated with classical approaches, redefining protein structure prediction problems from a high-dimensional perspective.
As protein sequences lengthen and the number of interactions increases, the corresponding Hamiltonian grows more complex, and the depth of the quantum circuit also increases. Consequently, noise accumulation becomes more pronounced, which can slow VQE convergence, trap the system in local minima, or shift the estimated energy away from its true value, all of which undermine the accuracy of the protein backbone conformation. Meanwhile, if the number of measurements in the final sampling stage is insufficient, statistical (shot) noise remains large; if measurements are too frequent, hardware is required to maintain the same circuit state for a longer duration, possibly amplifying noise further [35]. Some studies have demonstrated that using coarse-grained or lower-dimensional energy functions to represent protein structures can reduce the reliance on fully atomistic simulations and significantly decrease the sensitivity of quantum systems to noise [36, 37, 38]. Meanwhile, recent research provides further evidence that non-fault-tolerant quantum devices still possess sufficient performance to address utility tasks [39]. In our work, we employed a simplified vector-based protein backbone model, which has been shown to exhibit high robustness [40, 41], reducing the dimensionality of observables and the overall state space. Although this coarse-grained model cannot perfectly capture atomic-scale interactions, it is sufficient to represent key geometric features and energy trends. Our five model workflow also significantly reduces the impact of noise. In the first stage (estimation), the system avoids large-scale quantum-state measurements and focuses on rapidly estimating total energy, thereby reducing the accumulation of quantum gate and readout errors. In the second stage (sampling), after the circuit parameters have approximately converged, we perform extensive measurements in a single run to fully extract the final quantum-state distribution, applying readout error mitigation techniques to correct for systematic biases in the measurement process. While utility-level quantum devices still exhibit disturbances caused by noise, these disturbances do not have a practical impact on the results or significantly affect the overall performance of the system in the problem scenarios we addressed (Figures 2 and 4).
To mitigate noise in quantum processors, previous studies have proposed a range of techniques, such as adding error detection or error correction codes (ECC) to quantum circuits [42, 43, 44], or deploying denoisers at the pulse or measurement level [45, 46]. While these measures effectively reduce error rates, they also require additional physical qubits and quantum gate operations, increasing hardware demands and runtime [47]. Additionally, zero-noise extrapolation (ZNE) artificially amplifies circuit noise by inserting pairs of quantum gate operations, measures energy values under different noise levels, and extrapolates results toward a ”zero-noise” limit  [48]. Pulse-level optimizations and dynamic error correction may further suppress decoherence [49]. As hardware evolves and quantum error-correcting codes (QEC) mature, gate fidelities and the number of available qubits are expected to increase, providing broader possibilities for larger-scale and higher-accuracy protein structure prediction [50, 51, 52, 53]. Looking ahead, advancements in quantum hardware, including increased qubit numbers and improved error-correction capabilities, will enable quantum computing to address more complex biological tasks. Potential application areas include long-chain protein structure prediction, modeling multi-molecular interactions, and drug molecule design. Moreover, hybrid approaches combining classical optimizers with deep learning algorithms may further enhance the efficiency and applicability of quantum computing, driving revolutionary advancements in computational biology.
Methods
Data Source and Preparation
The data used in this study were obtained from the PDBbind database [11, 12, 13], which contains high-quality protein-ligand complex structures and protonation charge data. Protein structures with well-defined active sites were selected to ensure consistent structural information and charge states, thereby standardizing the inputs for the prediction models. Because our approach aims to provide a general solution for protein-ligand docking problems, rather than focusing on any specific drug discovery target, we randomly selected two sets of complexes from the PDBbind database for our experiments. There was no particular rationale for choosing these two sets beyond their availability and completeness; they serve as representative examples to demonstrate the universality of our method. The PDBbind database is widely accepted as a standard resource for evaluating docking and binding-affinity prediction methods and has been used in numerous previous studies [54, 55, 56, 57].
Geometric Encoding on a Tetrahedral Lattice
We adopted a tetrahedral lattice modeling approach inspired by previous studies [58] to construct the geometric backbone of proteins from amino acid sequences. Each amino acid node has four possible extension directions, represented by a set of directional vectors[59]. The distance between adjacent residues is fixed to unit length, and the bond angle is set to approximately 109.4°, consistent with fundamental protein geometry. This coarse-grained representation preserves the essential structural features of proteins while significantly reducing the number of required qubits. In our approach, the connection vectors between amino acid nodes, inter-residue interactions, and structural constraints are all mapped to a set of quantum operators. During computation, the number of qubits required by the quantum processor is determined by the dimensionality of these operators derived from the modeling process. In quantum systems, each qubit encodes a $2$-dimensional Hilbert space, so $n$ qubits can represent a space of dimension $2^{n}$. Once the system Hamiltonian is defined, the processor must allocate sufficient qubits to span the solution space. The number of qubits increases with problem complexity, which is jointly determined by the variety and length of the input amino acid sequence. Amino acid types influence the scale of interaction and constraint operators, while sequence length directly determines the dimensionality of the solution space. Table 1 summarizes the number of qubits and circuit depths used in each case. The core modeling logic exploits the superposition property of quantum systems. For example, if we encode an amino acid node $A$ using 4 qubits, each measurement outcome (e.g., 0001, 0010, 0100, or 1000) activates one of four extension vectors corresponding to directions in the tetrahedral lattice[60].. At the start, all possible connection paths for the sequence are encoded in superposition, where each amino acid node has four equally probable directions. Under the influence of the Hamiltonian, the variational optimization gradually alters the probability distribution of each path. During measurement, the quantum state collapses to the most probable configuration, represented by a binary string in which only one direction is activated per node. This outcome corresponds to the most likely amino acid connection determined by energy minimization. Since the Hamiltonian encodes constraints as energy penalties, invalid configurations lead to higher energy terms and are suppressed during optimization. As a result, the energy landscape of the protein is reflected in the energy domain of the quantum system, allowing quantum computation to efficiently search the conformational space via its inherent superposition capability. The quantum state evolves toward the conformation that minimizes the protein’s total energy.
Hamiltonian Design
The total Hamiltonian for protein structure prediction, $H_{\text{t}}$, consists of four main components:
$H_{\text{t}}=H_{\text{c}}+H_{\text{g}}+H_{\text{d}}+H_{\text{i}}$
$H_{\text{c}}$: Penalizes configurations that violate the natural chirality of amino acids, ensuring stereochemical correctness. $H_{\text{g}}$: Imposes geometric constraints to maintain backbone integrity. $H_{\text{d}}$: Penalizes unreasonable distance interactions, preventing spatial collisions. $H_{\text{i}}$: Encodes interaction energies between amino acids to identify the lowest-energy structure[61, 62, 63].
By minimizing $H_{\text{t}}$, the quantum system searches for the optimal protein conformation that satisfies both geometric and physical constraints. We utilized the Variational Quantum Eigensolver (VQE) to identify the ground state of the system and efficiently encoded the Hamiltonian using the SparsePauliOp class in Qiskit. The EfficientSU2 Ansatz was chosen for its balance between expressiveness and robustness against device noise, while the COBYLA algorithm served as the classical optimizer, suitable for utility-level environments.
We further introduced weighting factors $\lambda_{c},\lambda_{g},\lambda_{d},$ and $\lambda_{i}$ (all set to 1 in our current implementation) to adjust the relative strength of each term:
$H_{\text{t}}=\lambda_{c}H_{\text{c}}+\lambda_{g}H_{\text{g}}+\lambda_{d}H_{\text{d}}+\lambda_{i}H_{\text{i}}.$
These weights can be tuned to emphasize particular aspects of the protein structure, such as tighter chirality constraints or stronger inter-residue interactions. In our experiments, we found that setting all $\lambda$’s to the same value provided a reasonable balance for diverse protein sequences. However, specific cases may benefit from adjusting these coefficients, especially when dealing with atypical residues or post-translational modifications.
Quantum Circuit Design and Variational Optimization
In realistic scenarios, the qubit requirement grows significantly with the amino acid sequence length. For very short sequences, fewer than 10 qubits may suffice. However, in our experiments, when dealing with 12 amino acids, at least 77 qubits are needed to encode all relevant operators. Moreover, to satisfy certain coupling conditions and allow for additional SWAP operations that reduce overall circuit depth, we reserve about 5 extra qubits (totaling 82) as a buffer. This strategy effectively lowers the entangling-gate overhead (see, e.g., [64, 65] for related approaches). In many recent studies, the EfficientSU2 ansatz has emerged as a widely used and robust parameterized circuit, recognized for its relatively broad applicability and moderate resource requirements [66, 67, 68]. To parameterize our circuits, we employ the EfficientSU2 ansatz from Qiskit. This ansatz typically consists of repeated blocks of single-qubit rotation gates $R_{y}$ and $R_{z}$, interspersed with CX gates arranged in a circular or linear topology to introduce entanglement. By adjusting the number of repetition layers (reps), we can flexibly control the circuit’s expressive power. In a single repetition layer, each qubit undergoes one $R_{y}R_{z}$ rotation, followed by an entangling step that couples adjacent qubits. The theoretical circuit depth of Qiskit’s EfficientSU2 parameterized quantum circuit grows linearly with the number of qubits $N$. For our case with $N=82$ qubits: Total theoretical depth is $333$. The actual compiled depth may be different through gate optimization and hardware topology adaptation. An important advantage of the quantum method is its extremely fast computational speed. Due to the superposition state of the quantum system, during each optimization, the system does not focus on a specific structure for validation like traditional structure-based prediction methods. Instead, the superposition state allows all possible conformations of the current protein structure to exist simultaneously at the physical level, collapsing into a definite ground state as the optimal result through parameter optimization. In our tests, for a sequence with a length of 5 amino acids, the total recorded execution time after 150 iterations was 4079.55 seconds. According to logs from IBM’s superconducting quantum computer, the average time per optimization iteration and state measurement, from task submission to result acquisition, was approximately 10 seconds, with about 73.53% of the time spent on the quantum end. In our setup, the number of shots for optimization tasks was 200, and the number of shots for measurement was 100,000, indicating that the execution efficiency of the quantum computing method is extremely high. The quantum processor itself completes a single task almost instantaneously, with most of the time spent on state preparation and the machine’s self-calibration. Additionally, part of the time was limited by the performance of the classical computing backend. Under computational models of equivalent accuracy, the quantum method is several orders of magnitude faster than all current methods based on classical computing.
Docking Evaluation with AutoDock Vina
To evaluate the structural quality and potential functional effectiveness of the predicted protein fragment conformations, we conducted docking-based validation using AutoDock Vina. To ensure fairness and reproducibility, all structures—those predicted by our quantum computing framework and those generated by AlphaFold3 (AF3)—were subjected to a standardized preprocessing and docking pipeline as follows:
Structure preprocessing. All predicted structures were preprocessed using AutoDockTools before docking: polar hydrogens were added, Gasteiger charges were assigned, and backbone conformations were preserved. Receptor structures were derived from native PDBbind entries, and the ligands were held constant across all tests. The goal was to evaluate how different predicted conformations influence binding capability with the same ligand.
Docking protocol. For each test case, the ligand was docked to the predicted protein fragment using a grid box centered on the native binding site. Each structure underwent 20 independent docking runs, with each run producing 9 docking poses ranked by predicted binding energy, resulting in a total of 180 docking results per structure for statistical analysis.
Evaluation metrics. AutoDock Vina outputs were assessed using three key metrics: (1) Binding affinity score (kcal/mol), where lower scores indicate stronger binding potential; (2) RMSD lower bound (RLB), the minimum structural deviation between the top-scoring pose and all others, reflecting prediction stability; and (3) RMSD upper bound (RUB), the maximum deviation among all poses, indicating overall reliability[28, 29].
Summary of results. All docking results were analyzed statistically across the 20 repeated experiments. As shown in Figure 3B–D, structures predicted using the quantum method consistently outperformed those from AF3 in terms of both binding affinity and RMSD-based stability, demonstrating greater consistency and accuracy in geometric conformation and functional behavior.
Quantum Hardware and Execution Environment
All quantum computing tasks in this study were executed on real quantum hardware provided by IBM Quantum, specifically the IBM-Cleveland Clinic quantum processor. This device belongs to IBM’s Eagle r3 series within the IBM One line and features high-fidelity gate operations and a relatively stable readout mechanism, making it suitable for small- to medium-scale variational quantum algorithm experiments.
Task submission and management. All tasks were submitted via IBM’s Runtime service, using Qiskit’s dynamic scheduling capabilities to automatically allocate backend resources[69]. Each structure prediction task comprised two stages: Stage 1 executes the Variational Quantum Eigensolver (VQE) algorithm to minimize the ground-state energy; Stage 2 recompiles the circuit with fixed optimal parameters and performs large-scale measurements to extract the distribution of the final quantum state. The entire task flow was encapsulated into a two-stage session job and submitted to IBM’s job queue for execution.
Circuit design and compilation optimization. To align with the hardware topology of IBM-Cleveland Clinic, we applied gate remapping and gate fusion strategies in our quantum circuit design to minimize circuit depth and the number of two-qubit gates. Additionally, we utilized Qiskit’s transpiler stack to maximize execution fidelity while preserving circuit logic. For key tasks, noise mitigation techniques such as dynamic decoupling and readout error mitigation were employed to reduce systematic errors.
Execution time and resource consumption. Table 1 summarizes key statistics for each task, including sequence length, number of qubits used, circuit depth, optimization time, and energy range. Most tasks completed within acceptable time frames, with the most complex jobs requiring no more than 24 hours. It is worth noting that execution time can vary due to IBM’s backend scheduling algorithm: jobs that consume more resources may be placed in lower-priority queues and experience delays while awaiting higher-priority executions. These outliers reflect normal scheduling overheads. As sequence length increases, resource demands (qubit count and circuit depth) grow significantly; however, all tasks remained operationally stable on current utility-level devices[70].
Hardware limitations and mitigation strategies. Although today’s physical quantum devices are still limited by gate errors, readout noise, and finite coherence times, our two-stage scheduling mechanism and circuit compilation strategies successfully mitigated performance degradation from long-duration runs. This execution framework demonstrates the engineering feasibility of performing structure prediction on present-day utility-level quantum hardware.
Acknowledgments
Funding:
This work was primarily supported by the National Institute on Aging (NIA) under Award Numbers U01AG073323, R21AG083003, R01AG066707, R01AG076448, R01AG082118, R01AG084250, P30AG072959, and RF1AG082211, and the National Institute of Neurological Disorders and Stroke (NINDS) under Award Number RF1NS133812, and Alzheimer’s Association (ALZDISCOVERY-1051936) to F.C. This work was also supported in part by NIH grants U01 HG007691, R01 HL155107, and R01 HL166137; by AHA grants AHA957729 and AHA24MERIT1185447; and by EU Horizon 2021 grant 101057619 to J.L. This project has been funded in whole or in part with federal funds from the National Cancer Institute, National Institutes of Health, under contract HHSN261201500003I. The content of this publication does not necessarily reflect the views or policies of the Department of Health and Human Services, nor does mention of trade names, commercial products, or organizations imply endorsement by the U.S. Government. This Research was supported [in part] by the Intramural Research Program of the NIH, National Cancer Institute, Center for Cancer Research. This work was also partially supported by NSF 2238734 and NSF 2311950.
Author contributions:
F.C. and Q.G. conceived the study. Y.Z. designed research; Y.Z., Y.Y. and W.M. performed research; J.L., R.N., Y.Z., Y.Y., W.M., Z.W., C.L. and W.J. interpreted the data analysis; Y.Z., Y.Y., W.M., Q.G., F.C., K.L., R.N., and J.L. wrote and critically revised the manuscript. All authors gave final approval of the manuscript.
Competing interests:
There are no competing interests to declare.
Data and materials availability:
All data, code, and materials used in the analysis are available at Zenodo under the DOI: https://doi.org/10.5281/zenodo.15653560. This includes the full source code for the quantum structure prediction framework, Hamiltonian construction scripts, VQE configurations. The repository enables full reproduction of all results presented in this study.
Supplementary materials
Supplementary Text
Figs. S1 to S7
Supplementary Materials for
Prediction of Protein Three-dimensional Structures via a Hardware-Executable Quantum Computing Framework
Yuqi Zhang, Yuxin Yang, William Martin, Kingsten Lin, Zixu Wang, Cheng-Chang Lu, Weiwen Jiang, Ruth Nussinov, Joseph Loscalzo, Qiang Guan, Feixiong Cheng
Corresponding author. Email: qguan@kent.edu; chengf@ccf.org.
This PDF file includes:
Supplementary Text
Figures S1 to S7
Coloring Scheme for Protein Visualization in Figure4
In protein visualization, amino acids are categorized based on their chemical properties, and each category is assigned a specific color to enhance structural understanding. The coloring rules are as follows: 1. Hydrophobic Amino Acids: This group includes alanine (ALA), glycine (GLY), valine (VAL), leucine (LEU), isoleucine (ILE), and proline (PRO). These amino acids tend to be located in the interior of proteins to avoid contact with water. They are colored cyan to represent their hydrophobic nature. 2. Negatively Charged Amino Acids: This group includes aspartic acid (ASP) and glutamic acid (GLU). These amino acids have side chains that carry a negative charge at physiological pH, making them acidic. They are colored yellow to highlight their negative charge. 3. Positively Charged Amino Acids: This group includes arginine (ARG), lysine (LYS), and histidine (HIS). These amino acids have side chains that carry a positive charge at physiological pH, making them basic. They are colored red to represent their positive charge. 4. Polar Uncharged Amino Acids: This group includes serine (SER), threonine (THR), asparagine (ASN), and glutamine (GLN). These amino acids have polar side chains but do not carry a charge at physiological pH. They are colored green to reflect their polarity. 5. Sulfur-Containing Amino Acids: This group includes cysteine (CYS) and methionine (MET). These amino acids contain sulfur atoms and play crucial roles in protein function and structural stability. They are colored magenta to emphasize their unique chemical properties. 6. Aromatic Amino Acids: This group includes phenylalanine (PHE), tyrosine (TYR), and tryptophan (TRP). These amino acids have conjugated aromatic ring structures, which are important for protein stability and interactions. They are colored orange to highlight their aromatic nature. This coloring scheme is designed to visually distinguish the chemical properties and spatial distribution of amino acids in a protein structure, aiding in the interpretation of protein function and characteristics.
Figure S1: RMSD comparison between quantum and AF3 predictions. In 18 out of 23 test cases, quantum-predicted structures exhibit lower RMSD values relative to experimental structures compared to AF3, indicating improved geometric fidelity.
Figure S2: Docking affinity scores of quantum vs. AF3-predicted structures. Quantum structures achieved more favorable (lower) binding free energies in 21 of 23 cases, indicating better predicted protein-ligand interactions than AF3.
Figure S3: Parameterized quantum circuit design used in VQE. The ansatz follows the EfficientSU2 structure with alternating entanglement and rotation layers, enabling expressive quantum state preparation for structure optimization.
Figure S4: Predicted structures and energy profiles for 1a9m, 1fkn, and 1qin. Energy trajectories during VQE optimization are shown alongside the corresponding low-energy conformations for each fragment.
Figure S5: Predicted structures and energy profiles for 3ans and 3b26. Conformations sampled during VQE convergence illustrate how quantum-optimized structures correspond to energy minima.
Figure S6: Coupling map of the IBM-Cleveland Clinic quantum processor. The hardware topology used in this study supports 127 qubits with directional two-qubit gate connectivity optimized for low error rates.
Figure S7: Raw noise statistics distribution of the IBM-Cleveland Clinic quantum processor.